# Counting rational points on smooth quartic and quintic surfaces

### v1

@@ Abstract.
Let $X$ be a smooth projective surface of degree $d\geq 4$ defined over a number field $K$, and let $N_{X^{\prime}}(B)$ be the number of rational points of $X$ of height at most $B$ that do not lie on lines contained in $X$. Assuming a suitable hypothesis on the size of the rank of abelian varieties, we show that $N_{X^{\prime}}(B)\ll_{X}B^{4/3+\varepsilon}$ for any fixed $\varepsilon>0$. This improves an unconditional and uniform bound from Salberger for $d=4$ and $d=5$. The proof, based on an argument of Heath-Brown, consists of cutting $X$ by projective planes and using a uniform version of Faltings’s Theorem due to Dimitrov, Gao, and Habegger, to bound the number of rational points on the plane sections of $X$.
@@ 1. Introduction
$H_{K}([x_{0}\colon\cdots\colon x_{n}])=\prod_{v}\max\{\lVert x_{1}\rVert_{v},\dotsc\lVert x_{n}\rVert_{v}\},$
where $v$ runs through the places of $K$ and $\lVert x\rVert_{v}$ is the normalized absolute value associated with the place $v$ (this is the height function defined in [hindry, Section B.2]).
We consider a smooth hypersurface $X\subseteq\mathbb{P}_{K}^{n}$ of degree $d$, and we want to bound the number of rational points of bounded height on $X$. For this, define the counting function
@@
For $K=\operatorname{\mathbb{Q}}$, Verzobio shows in [verzobio] that for $n\geq 4$ and $d\geq 50$ we have
$N_{X}(B)\ll_{n,d,\epsilon}B^{n-2+\varepsilon},$
@@
For $n=3$ and $K=\operatorname{\mathbb{Q}}$, Salberger shows in [salberger, Theorem 0.5] the bound
@@
We aim to improve this bound (although not uniformly in $X$) for $d=4$ and $5$, using a technique developed by Heath-Brown in [heath-brown] and assuming a conjectural bound for the rank of abelian varieties. In his paper, Heath-Brown shows that for $d=3$ and assuming such a hypothesis, we have, for any fixed $\varepsilon>0$, a bound
@@
We will cover the points of bounded height with projective planes of bounded height in the dual projective space $(\mathbb{P}^{3})^{\vee}$. This can be done using a result from Schmidt ([schmidt]). We then apply another result from the same paper to bound the number of such planes asymptotically, and proceed by estimating the number of points of bounded height in the intersection between $X$ and each of these planes. This intersection will be generically a smooth curve. In this case, we use a uniform version of Faltings’s Theorem, due to Dimitrov, Gao and Habegger:
Lemma 1.1 (Uniform Faltings’s Theorem ([mordell-lang])).
For each integer $g\geq 2$, there exists a constant $c(g)>0$ depending on $K\colon\operatorname{\mathbb{Q}}$] such that, for every curve $C$ over $K$ of (geometric) genus $g$, we have
@@
This is [mordell-lang, Theorem 1]. Compare it also with [kuhne, Theorem 4], which eliminates the dependency on the degree $K:\operatorname{\mathbb{Q}}]$.
@@
We can proceed similarly to bound the number of rational points on the irreducible components of genus $1$, by using a uniform result due to Heath-Brown and Testa ([hb-testa]). Finally, we show that the curves of genus 0 that appear do not alter our estimation.
@@
This hypothesis is formulated in [heath-brown] for elliptic curves over the rational numbers.
We remark that what we actually need is the a priori weaker bound $r_{A}=o(h_{F}(A))$, where $h_{F}(A)$ denotes the Faltings’s height of $A$. However, the plausibility of the Rank Hypothesis is suggested by the following analytical argument:
@@
which gives the Rank Hypothesis. See [mestre, Proposition II.1] for a detailed proof of these implications in the case of the $L$-function of a modular form.
We now cite some known results that go in the direction of the Rank Hypothesis. By [ooe, Theorem 1], for an abelian variety $A/K$ we have, unconditionally,
@@
In the case of elliptic curves, more evidence is known. Notice that all elliptic curves over the rational number are modular by [breuil]. If we assume that an elliptic curve $E/\operatorname{\mathbb{Q}}$ has at least one rational point of order $2$, then it is known that (see [heath-brown, Section 6])
@@
Assume that the Rank Hypothesis holds. Let $X\subseteq\mathbb{P}^{3}$ be a smooth surface of degree $d\geq 4$, and let $\varepsilon>0$. Then $N_{X^{\prime}}(B)\ll_{X,K,\varepsilon}B^{4/3+\varepsilon}$.
This exponent is smaller than the one in (1) for $d=4$ and $d=5$, thus improving the known bound in this case.
[…]
We remark that, assuming the stronger conjecture that the rank of all abelian varieties is bounded in terms of their dimensions and the field $K$, our proof will show that we can remove the dependence on $\varepsilon$ and obtain
$N_{X^{\prime}}(B)\ll_{X,K}B^{4/3}.$
Some of the references use a slightly different definition of $H_{K}$, taking the Archimedean norm in infinite places instead of the maximum norm. However, since both norms differ by a multiplicative constant, this distinction will not matter in what follows, so we shall ignore it.
By genus, we always mean geometric genus, unless explicitly stated. Besides, we will denote by $x_{0},x_{1},x_{2},x_{3}$ the coordinates of $\mathbb{P}^{3}$ and by $F(x_{0},x_{1},x_{2},x_{3})$ the form of degree $d$ that defines $X$.
@@
Let $V$ be an $n$-dimensional $K$-vector space, and let $v\in V$. Let $H_{K}$ be the height on $V$ given by the identification $V=\operatorname{\mathbb{A}}_{K}^{n}(K)\xhookrightarrow{}\mathbb{P}_{K}^{n}(K)$. Let $1\leq k\leq n-1$. We start by defining the height in the Grassmannian
@@
In our argument, we will only use $\operatorname{\operatorname{Gr}}_{K}(3,4)$, which is the dual projective space $(\mathbb{P}^{3})^{\vee}$. However, the more general concept of a Grassmanian will be useful if one wants to generalize this argument in higher dimensions.
Let $W\subseteq V$ be a $k$-dimensional subspace. Let $w_{1},\dotsc,w_{k}$ be any basis of $W$. Consider the exterior product $w_{1}\wedge\cdots\wedge w_{k}\in\bigwedge^{k}V\cong K^{\binom{n}{k}}$. We define $H_{K}(W)\coloneqq H_{K}(w_{1}\wedge\cdots\wedge w_{k})$. It is possible to show that this height does not depend on the choice of basis, so $H_{K}(W)$ is well-defined. For more on the heights of subspaces, see [schmidt].
Returning to our original problem, let $x\in X(K)$. We want to find a projective plane $\Pi\cong\mathbb{P}^{2}$ of small height containing $x$. This is equivalent to finding a $3$-dimensional subspace of a $4$-dimensional space that contains a given vector $x$. By [schmidt, Theorem 2], there exists such a plane $\Pi$ with
$H_{K}(\Pi)\ll_{K,n,k}H_{K}(x)^{\frac{4-3}{4-1}}=H_{K}(x)^{\frac{1}{3}}.$
This shows that, for a given $B>0$, the set
@@
can be covered by planes of height $O_{K}(B^{\frac{1}{3}})$. Using now [schmidt, Theorem 3] (or the more precise version [thunder, Theorem 1]), we see that the number of such planes is $O_{K}\big(\big(B^{\frac{1}{3}}\big)^{4}\big)=O_{K}(B^{4/3})$.
@@
Let $C$ be an irreducible degree $e\leq d$ component of this curve, and let $\tilde{C}$ be its normalization. We can compute the genus of $C$ as being
$g(C)=\frac{(e-1)(e-2)}{2}-\sum_{p}\frac{m_{p}(m_{p}-1)}{2},$ (2)
where the sum is taken over all singular points of $C$, including infinitely near singular points, and $m_{P}$ denotes the multiplicity of $C$ at $p$ (see, for example, [hartshorne, Example 3.9.2]). In particular, as $g(C)\geq 0$, we see that $C$ has a finite number of singular points, and this number is bounded in terms of $d$. Since $\tilde{C}$ coincides with $C$ outside the finite set of singular points, we see that counting the rational points of $C(K)$ and $\tilde{C}(K)$ is the same up to $O_{d}(1)$.
If $g(C)\geq 2$, we can apply the Uniform Faltings’s Theorem (Lemma 1.1) to $\tilde{C}$ to find
@@
$\tilde{C}(K)\ll_{X,K,\varepsilon}B^{\varepsilon}.$
Since $\tilde{C}$ is obtained by normalizing an irreducible component of the intersection $X\cap\Pi$ and $H_{K}(\Pi)\ll B^{1/3}$, the Jacobian $J$ of $\tilde{C}$ is defined by equations with coefficients of size $O_{X}(B^{1/3})$, thus
$h_{F}(J)\ll_{X}\log B,$
where $h_{F}(J)$ is the Faltings’s height. It is well-known that $\log N_{J}\ll h_{F}(J)$ (one can deduce this inequality for Abelian varieties with semistable reduction over $K$ by using [de-jong-shokrieh, Theorem A], and from this deduce the general case by using [hindry-pacheco, Lemma 3.4]). Therefore, by the Rank Hypothesis, we have
$\rho(\tilde{C})=r_{J}=o(\log N_{J})=o(h_{F}(J))=o(\log B).$

### v2

@@ Abstract.
Let $X\subseteq\mathbb{P}^{3}$ be a smooth projective surface of degree $d\geq 4$ defined over a number field $K$, and let $N_{X^{\prime}}(B)$ be the number of rational points of $X$ of height at most $B$ that do not lie on lines contained in $X$. Assuming a suitable hypothesis on the size of the rank of Abelian varieties, we show that $N_{X^{\prime}}(B)\ll_{K,d,\varepsilon}B^{4/3+\varepsilon}$ for any fixed $\varepsilon>0$. This improves an unconditional bound from Salberger for $d=4$ and $d=5$. The proof, based on an argument of Heath-Brown, consists of cutting $X$ by projective planes and using a uniform version of Faltings’s Theorem, due to Dimitrov, Gao, and Habegger, to bound the number of rational points on the plane sections of $X$. More generally, we prove that if $X\subseteq\mathbb{P}^{n}$ is a non-degenerate non-uniruled smooth projective surface defined over $K$, then $N_{X^{\prime}}(B)\ll_{K,n,d,\varepsilon}B^{\frac{n+1}{n}+\varepsilon}$.
@@ 1. Introduction
$H_{K}([x_{0}\colon\cdots\colon x_{n}])=\prod_{v}\max\{\lVert x_{0}\rVert_{v},\dotsc\lVert x_{n}\rVert_{v}\},$
where $v$ runs through the places of $K$ and $\lVert x\rVert_{v}$ is the normalized absolute value associated with the place $v$ (this is the height function defined in [10, Section B.2]).
We consider a smooth hypersurface $X\subseteq\mathbb{P}^{n}$ of degree $d$, and we want to bound the number of $K$-rational points of bounded height on $X$. For this, define the counting function
@@
For $K=\operatorname{\mathbb{Q}}$, Verzobio shows in [21] that for $n\geq 4$ and $d\geq 50$ we have
$N_{X}(B)\ll_{n,d,\varepsilon}B^{n-2+\varepsilon},$
@@
For $n=3$ and $K=\operatorname{\mathbb{Q}}$, Salberger shows in [18, Theorem 0.5] the bound
@@
We aim to improve this bound for $d=4$ and $5$, using a technique developed by Heath-Brown in [8] and assuming a conjectural bound for the rank of Abelian varieties. In his paper, Heath-Brown shows that for $d=3$ and assuming such a hypothesis, we have, for any fixed $\varepsilon>0$, a bound
@@
This technique was also applied by Glas and Hochfilzer to bound the number of rational points on Del Pezzo surfaces of low degree over a global field, assuming the validity of the same conjectural bound for number fields ([6]).
We will cover the points of bounded height with projective planes of bounded height in the dual projective space $(\mathbb{P}^{3})^{\vee}$. This can be done using a result from Schmidt ([19]). We then apply another result from the same paper to bound the number of such planes asymptotically, and proceed by estimating the number of points of bounded height in the intersection between $X$ and each of these planes. This intersection will be generically a smooth curve. In this case, we use a uniform version of Faltings’s Theorem, due to Dimitrov, Gao, and Habegger:
Lemma 1.1 (Uniform Faltings’s Theorem ([5])).
For each integer $g\geq 2$, there exists a constant $c(g)>0$ depending on $K\colon\operatorname{\mathbb{Q}}$] such that, for every smooth projective curve $C$ over $K$ of (geometric) genus $g$, we have
@@
This is [5, Theorem 1]. Compare it also with [12, Theorem 4], which eliminates the dependence on the degree $K:\operatorname{\mathbb{Q}}]$.
@@
We can proceed similarly to bound the number of rational points on the irreducible components of genus $1$, by using a bound on the number of smooth cubic plane curves due to Heath-Brown ([8]), which is generalized to global fields by Glas and Hochfilzer ([6]). Finally, we show that the curves of genus 0 that appear do not alter our estimation. We remove dependence on $X$ by applying a standard trick due to Heath-Brown.
@@
This hypothesis is formulated in [8] for elliptic curves over the rational numbers.
We remark that what we actually need is the a priori weaker bound $r_{A}=o(h_{\operatorname{Falt}}(A))$, where $h_{\operatorname{Falt}}(A)$ denotes the Faltings’s height of $A$. However, the plausibility of the Rank Hypothesis is suggested by the following analytical argument:
@@
which gives the Rank Hypothesis. See [13, Proposition II.1] for a detailed proof of these implications in the case of the $L$-function of a modular form.
We now cite some known results that go in the direction of the Rank Hypothesis. By [14, Theorem 1], for an Abelian variety $A/K$ we have, unconditionally,
@@
In the case of elliptic curves, more evidence is known. Notice that all elliptic curves over the rational number are modular by [1]. If we assume that an elliptic curve $E/\operatorname{\mathbb{Q}}$ has at least one rational point of order $2$, then it is known that (see [8, Section 6])
@@
Assume that the Rank Hypothesis holds. Let $X\subseteq\mathbb{P}^{3}$ be a smooth surface of degree $d\geq 4$, and let $\varepsilon>0$. Then we have $N_{X^{\prime}}(B)\ll_{K,d,\varepsilon}B^{4/3+\varepsilon}$.
This exponent is smaller than the one in (1) for $d=4$ and $d=5$, thereby improving the known bound in these cases.
Another advantage of our method is that it works over any number field, so we obtain the same bound for every choice of $K$. To do this, we use results of Paredes and Sasyk ([15]) and Glas and Hochfilzer ([6]). However, notice that we use the height $H_{K}$ of the field $K$, while in [15] they use the absolute height $H=H_{K}^{1/[K:\operatorname{\mathbb{Q}}]}$.
We can generalize our result for smooth surfaces contained in some bigger projective space:
Theorem 1.4.
Assume that the Rank Hypothesis holds. Let $n\geq 3$, let $X\subseteq\mathbb{P}^{n}$ be a non-degenerate non-uniruled smooth surface of degree $d$, and let $\varepsilon>0$. Then we have $N_{X^{\prime}}(B)\ll_{K,n,d,\varepsilon}B^{\frac{n+1}{n}+\varepsilon}$.
The proof of this result is analogous to the proof of Theorem 1.3, except that we cover $X$ by hyperplanes of $\mathbb{P}^{n}$ instead of by planes of $\mathbb{P}^{3}$. Since a uniruled surface always has Kodaira dimension -$\infty$, the above theorem applies to surfaces of general type, general elliptic surfaces, abelian surfaces, hyperelliptic surfaces, Enriques surfaces and K3 surfaces.
We hope to generalize this method in a future paper to obtain results about higher dimensional varieties.
Some articles use a slightly different definition of $H_{K}$, taking the Archimedean norm in infinite places instead of the maximum norm. However, since both norms differ by a multiplicative constant, this distinction will not matter in what follows, so we shall ignore it.
Given a polynomial $P(x_{0},\dotsc,x_{n})\in K[x_{0},\dotsc,x_{n}]$, we define its height $H_{K}(P)$ as the height over $K$ of the projective point whose entries are the coefficients of $P$, and its logarithmic height to be $h_{K}(P)\coloneqq\log H_{K}(P)$.
By genus, we always mean the geometric genus, unless explicitly stated. Besides, we will denote by $x_{0},x_{1},x_{2},x_{3}$ the coordinates of $\mathbb{P}^{3}$ and by $F(x_{0},x_{1},x_{2},x_{3})$ the form of degree $d$ that defines $X$.
Following [6], we define, for $x=(x_{0},\dotsc,x_{n})\in\mathcal{O}_{K}^{n+1}$:
$\lVert x\rVert\coloneqq\max_{v\mid\infty}\max\{\lVert x_{0}\rVert_{v},\dotsc,\lVert x_{n}\rVert_{v}\},$
and
$\lVert x\rVert_{\infty}\coloneqq\prod_{v\mid\infty}\max\{\lVert x_{0}\rVert_{v},\dotsc,\lVert x_{n}\rVert_{v}\}.$
Fix $\mathfrak{a}_{1},\dotsc,\mathfrak{a}_{h}\subseteq\mathcal{O}_{K}$ to be a full set of representatives in the class group of $K$. Whenever we indicate that an implied constant depends on $K$, we implicitly allow it to also depend on our choice of the ideal class group representatives. Define
$Z_{n}^{\prime}\coloneqq\{x\in\mathcal{O}_{K}^{n+1}\setminus\{0\}\colon(x_{0},\dotsc,x_{n+1})=\mathfrak{a}_{i}\text{ for some }i=1,\dotsc,h\}.$
Let $s_{K}$ be the number of infinite places of $K$. By [6, Lemma 2.8], there exist constants $c_{1},c_{2}>0$ depending on $K$ such that every member of $\mathbb{P}^{n}(K)$ has a representative $x\in Z_{n}^{\prime}$ such that
$c_{1}\lVert x\rVert\leq\lVert x\rVert_{\infty}^{1/s_{K}}\leq c_{2}\lVert x\rVert.$
Hence, if we define
$Z_{n}\coloneqq\{x\in Z_{n}^{\prime}\colon c_{1}\lVert x\rVert\leq\lVert x\rVert_{\infty}^{1/s_{K}}\leq c_{2}\lVert x\rVert\},$
then every point of $\mathbb{P}^{n}(K)$ has a representative in $Z_{n}$. Notice that, for $x\in Z_{n}^{\prime}$, $H_{K}(x)$ is of the same order as $\lVert x\rVert_{\infty}$, since $H_{K}(x)$ is equal to $\lVert x\rVert_{\infty}$ divided by the norm of the ideal $(x_{0},\dotsc,x_{n+1})=\mathfrak{a}_{i}$ for some $1\leq i\leq h$.
@@
We start with the following dichotomy:
Lemma 2.1.
Either $H_{K}(F)\ll_{K}B^{d\binom{d+3}{3}}$ or there exists a surface $Y\subseteq\mathbb{P}^{3}$ of degree $d$ such that $X\cap Y$ is a curve, and all $K$-rational points of $X$ of height smaller than $B$ are contained in this curve.
Proof.
This is a generalization of [3, Lemma 5] to number fields. We can assume that $F$, the form defining $X$, has coefficients in $Z_{\binom{d+3}{3}-1}$. Suppose that $x^{(1)},\dotsc,x^{(N)}$ are all the rational points of $X$ satisfying $H_{K}(x)\leq B$, and write these points with representatives in $Z_{3}$. Let $M$ be the $\binom{d+3}{3}\times N$ matrix whose $i$-th row consists of the $\binom{d+3}{3}$ possible monomials of degree $d$ in the variables $x_{0}^{(i)},\dotsc,x_{3}^{(i)}$.
If $f$ is the vector given by the coefficients of $F$, then $Mf=0$, which shows that the rank of $M$ is at most $\binom{d+3}{3}-1$. Thus, we can find a nonzero vector $g\in Z_{\binom{d+3}{3}-1}$ constructed from subdeterminants of $M$ such that $Mg=0$. Hence $\lVert g\rVert\ll B^{d\binom{d+3}{d}/s_{K}}$. Let $G$ be the form of degree $d$ corresponding to the vector $g$, and $Y\subseteq\mathbb{P}^{3}$ be the surface defined by $G$. Then $G(x^{(i)})=0$ for all $1\leq i\leq N$, hence $\{x^{(1)},\dotsc,x^{(N)}\}\subseteq X\cap Y$. If $G$ is not proportional to $F$, we are done. Otherwise, since both $f$ and $g$ are in $Z_{\binom{d+3}{3}-1}$, we have
$\lVert f\rVert\ll\lVert g\rVert\ll B^{d\binom{d+3}{3}/s_{K}}.$
Finally, we conclude that
$H_{K}(F)=H_{K}(f)\ll_{K}\lVert f\rVert_{\infty}\ll_{K}\lVert f\rVert^{s_{K}}\ll B^{d\binom{d+3}{3}}$.
If we are in the second case of the lemma above, then $N_{X^{\prime}}(B)\leq N_{(X\cap Y)^{\prime}}(B)$, where $X\cap Y$ is a curve. This curve has a finite number of components. Let $C\subseteq X\cap Y$ be an irreducible component of degree $e\geq 2$. By [15, Theorem 1.8],
$N_{C}(B)\ll_{K,d,\varepsilon}B^{2/e+\varepsilon}\leq B^{1+\varepsilon}.$
If $e=1$, then $C$ is a line, and by definition it does not contribute to the calculation of $N_{X^{\prime}}(B)$. In this case, we conclude that $N_{X^{\prime}}\ll_{K,d,\varepsilon}B^{1+\varepsilon}<B^{4/3+\varepsilon}$. Therefore, from now on, we will assume that we are in the first case of the lemma above, that is, $H_{K}(F)\ll B^{d\binom{d+3}{3}}$.
Let $V$ be an $n$-dimensional $K$-vector space, and let $v\in V$. Let $H_{K}$ be the height on $V$ given by the identification $V=\operatorname{\mathbb{A}}_{K}^{n}(K)\xhookrightarrow{}\mathbb{P}^{n}(K)$. Let $1\leq k\leq n-1$. We start by defining the height in the Grassmannian
@@
In our argument, we will only use $\operatorname{\operatorname{Gr}}_{K}(3,4)$, which is the dual projective space $(\mathbb{P}^{3})^{\vee}$. However, the more general concept of a Grassmanian will be useful if one wants to generalize this argument for varieties of higher dimension.
Let $W\subseteq V$ be a $k$-dimensional subspace. Let $w_{1},\dotsc,w_{k}$ be any basis of $W$. Consider the exterior product $w_{1}\wedge\cdots\wedge w_{k}\in\bigwedge^{k}V\cong K^{\binom{n}{k}}$. We define $H_{K}(W)\coloneqq H_{K}(w_{1}\wedge\cdots\wedge w_{k})$. It is possible to show that this height does not depend on the choice of basis, so $H_{K}(W)$ is well-defined. For more on the heights of subspaces, see [19].
Returning to our original problem, let $x\in X(K)$. We want to find a projective plane $\Pi\cong\mathbb{P}^{2}$ of small height containing $x$. This is equivalent to finding a $3$-dimensional subspace of a $4$-dimensional space that contains a given vector $x$. By [19, Theorem 2], there exists such a plane $\Pi$ with $H_{K}(\Pi)\ll_{K}B^{1/3}$. This shows that, for a given $B>0$, the set
@@
can be covered by planes of height $O_{K}(B^{1/3})$. Using now [19, Theorem 3] (or the more precise version [20, Theorem 1]), we see that the number of such planes is
$O_{K}\big(\big(B^{1/3}\big)^{4}\big)=O_{K}(B^{4/3}).$
@@
Let $C$ be an irreducible degree $e\leq d$ component of this curve, and let $\tilde{C}$ be its normalization. We can compute the genus of $\tilde{C}$ as being
$g(\tilde{C})=\frac{(e-1)(e-2)}{2}-\sum_{p}\frac{m_{p}(m_{p}-1)}{2},$ (2)
where the sum is taken over all singular points of $C$, including infinitely near singular points, and $m_{P}$ denotes the multiplicity of $C$ at $p$ (see, for example, [7, Example 3.9.2]). In particular, as $g(\tilde{C})\geq 0$, we see that $C$ has a finite number of singular points, and this number is bounded in terms of $d$. Since $\tilde{C}$ coincides with $C$ outside the finite set of singular points, we see that counting the rational points of $C(K)$ and $\tilde{C}(K)$ is the same up to $O_{d}(1)$.
If $g=g(\tilde{C})\geq 2$, we can apply the Uniform Faltings’s Theorem (Lemma 1.1) to $\tilde{C}$ to find
@@
$\tilde{C}(K)\ll_{K,d,\varepsilon}B^{\varepsilon}.$
We start by observing that, since $g\leq\frac{(d-1)(d-2)}{2}$, any dependence on $g$ can be transformed into a dependence on $d$. We will use this fact several times in what follows.
The curve $\tilde{C}$ is obtained by normalizing an irreducible component of the intersection $X\cap\Pi$, where $H_{K}(\Pi)\ll_{K}B^{1/3}$. This means that we have an equation for $\Pi$ of the form
$a_{0}x_{0}-a_{1}x_{1}-a_{2}x_{2}-a_{3}x_{3}=0,$
where $H_{K}([a_{0}\colon a_{1}\colon a_{2}\colon a_{3}])\ll_{K}B^{1/3}$. Supposing without loss of generality $a_{0}\neq 0$, the intersection $X\cap\Pi$ is defined by the equation:
$F_{\Pi}(x_{1},x_{2},x_{3})\coloneqq F(a_{1}x_{1}+a_{2}x_{2}+a_{3}x_{3},a_{0}x_{1},a_{0}x_{2},a_{0}x_{3})=0.$
Write
$F(x_{0},x_{1},x_{2},x_{3})$ $=$ $\sum_{\begin{subarray}{c}e\in\operatorname{\mathbb{N}}^{4},\\ \lvert e\rvert=d\end{subarray}}b_{e}x_{0}^{e_{0}}x_{1}^{e_{1}}x_{2}^{e_{2}}x_{3}^{e_{3}},\text{ and}$
$F_{\Pi}(x_{1},x_{2},x_{3})$ $=$ $\sum_{\begin{subarray}{c}f\in\operatorname{\mathbb{N}}^{3},\\ \lvert f\rvert=d\end{subarray}}c_{f}x_{1}^{f_{1}}x_{2}^{f_{2}}x_{3}^{f_{3}}.$
Expanding the expression for $F_{\Pi}$, we obtain:
$F_{\Pi}(x_{1},x_{2},x_{3})$ $=$ $\sum_{\begin{subarray}{c}e\in\operatorname{\mathbb{N}}^{4},\\ \lvert e\rvert=d\end{subarray}}b_{e}(a_{1}x_{1}+a_{2}x_{2}+a_{3}x_{3})^{e_{0}}x_{1}^{e_{1}}x_{2}^{e_{2}}x_{3}^{e_{3}}$
$=$ $\sum_{\begin{subarray}{c}e\in\operatorname{\mathbb{N}}^{4},\\ \lvert e\rvert=d\end{subarray}}b_{e}\sum_{\begin{subarray}{c}j\in\operatorname{\mathbb{N}}^{3},\\ \lvert j\rvert=e_{0}\end{subarray}}\binom{e_{0}}{j_{1},j_{2},j_{3}}(a_{1}x_{1})^{j_{1}}(a_{2}x_{2})^{j_{2}}(a_{3}x_{3})^{j_{2}}(a_{0}x_{1})^{e_{1}}(a_{0}x_{2})^{e_{2}}(a_{0}x_{3})^{e_{3}}$
$=$ $\sum_{\begin{subarray}{c}e\in\operatorname{\mathbb{N}}^{4},\\ \lvert e\rvert=d\end{subarray}}b_{e}\sum_{\begin{subarray}{c}j\in\operatorname{\mathbb{N}}^{3},\\ \lvert j\rvert=e_{0}\end{subarray}}\binom{e_{0}}{j_{1},j_{2},j_{3}}a_{0}^{e_{1}+e_{2}+e_{3}}a_{1}^{j_{1}}a_{2}^{j_{2}}a_{3}^{j_{3}}x_{1}^{e_{1}+j_{1}}x_{2}^{e_{2}+j_{2}}x_{3}^{e_{3}+j_{3}}$
$=$ $\sum_{\begin{subarray}{c}f\in\operatorname{\mathbb{N}}^{3},\\ \lvert f\rvert=d\end{subarray}}b_{(j_{1}+j_{2}+j_{3},f_{1}-j_{1},f_{2}-j_{2},f_{3}-j_{3})}\sum_{\begin{subarray}{c}j\in\operatorname{\mathbb{N}}^{3},\\ \lvert j\rvert\leq d\end{subarray}}\binom{j_{1}+j_{2}+j_{3}}{j_{1},j_{2},j_{3}}a_{0}^{d-j_{1}-j_{2}-j_{3}}a_{1}^{j_{1}}a_{2}^{j_{2}}a_{3}^{j_{3}}x_{1}^{f_{1}}x_{2}^{f_{2}}x_{3}^{f_{3}}.$
This shows that, for each $f\in\operatorname{\mathbb{N}}^{4}$, $\lvert f\rvert=d$, we have:
$c_{f}=b_{(j_{1}+j_{2}+j_{3},f_{1}-j_{1},f_{2}-j_{2},f_{3}-j_{3})}\sum_{\begin{subarray}{c}j\in\operatorname{\mathbb{N}}^{3},\\ \lvert j\rvert\leq d\end{subarray}}\binom{j_{1}+j_{2}+j_{3}}{j_{1},j_{2},j_{3}}a_{0}^{d-j_{1}-j_{2}-j_{3}}a_{1}^{j_{1}}a_{2}^{j_{2}}a_{3}^{j_{3}}.$
Therefore, for any place $v$ of $K$:
$\max_{\begin{subarray}{c}f\in\operatorname{\mathbb{N}}^{3},\\ \lvert f\rvert=d\end{subarray}}\lVert c_{f}\rVert_{v}\ll_{d}\max_{\begin{subarray}{c}e\in\operatorname{\mathbb{N}}^{4},\\ \lvert e\rvert=d\end{subarray}}\lVert b_{e}\rVert_{v}\cdot\big(\max_{0\leq i\leq 3}\lVert a_{i}\rVert_{v}\big)^{d}.$
Furthermore, if $v$ is a non-archimedean place, the explicit constant in the inequality above can be taken to be $1$. Multiplying the equation above for all $v$, we obtain:
$H_{K}(F_{\Pi})\ll_{d}H_{K}(F)\cdot H_{K}([a_{0}\colon a_{1}\colon a_{2}\colon a_{3}])^{d}\ll_{K}H_{K}(F)\cdot B^{d/3}.$
Taking the logarithm on both sides, we obtain an inequality of logarithmic heights:
$h_{K}(F_{\Pi})\ll_{K,d}h_{K}(F)+\log B+1.$
Hence, denoting by h(C~)h($\tilde${C}) the naive height of the curve $\tilde{C}$ (that is, the minimum logarithmic height of equations defining $\tilde{C}$), we have:
$h(\tilde{C})\leq h_{K}(F_{\Pi})\ll_{K,d}h_{K}(F)+\log B+1.$
Let $h_{\operatorname{Falt}}(J)$ be the Faltings height of $J$. We want to compare $h_{\operatorname{Falt}}(J)$ with h(C~)h($\tilde${C}). For this, we will also use $h_{\Theta}(J)$, the theta height of $J$ relative to the sixteenth power of a symmetric translation of the theta divisor of $J$. By [16, Corollary 1.3]:
$\lvert h_{\Theta}(J)-\frac{1}{2}\cdot h_{\operatorname{Falt}}(J)\rvert\ll_{g}\log(2+\max\{h_{\Theta}(J),1\}).$
This gives us an inequality:
$h_{\operatorname{Falt}}(J)\ll_{d}\max\{h_{\Theta}(J),1\}$
Using [17, Théorème 1.3 and Proposition 1.1] and using that $g$ is bounded in function of $d$, we obtain:
$h_{\Theta}(J)\ll_{d}h(\tilde{C})+1.$
Combining the two inequalities above and using the inequality obtained for the naive height gives:
$h_{\operatorname{Falt}}(J)\ll_{d}h(\tilde{C})+1\ll_{K,d}h_{K}(F)+\log B+1.$
It is well-known that $\log N_{J}\ll h_{\operatorname{Falt}}(J)$ (one can deduce this inequality for Abelian varieties with semistable reduction over $K$ by using [4, Theorem A], and from this deduce the general case by using [9, Lemma 3.4]). Therefore, by the Rank Hypothesis:
$\rho(\tilde{C})=r_{J}=o(\log N_{J})=o(h_{\operatorname{Falt}}(J))=o_{K,d}(h_{K}(F)+\log B).$